# Deep Context-Conditioned Anomaly Detection for Tabular Data

### v1

@@ 5.1.3. Loss Function
(1) $\quad+\lambda\cdot\text{MMD}\Big{(}P(z),Q(z\mid X=\mathbf{Y},\mathbf{C})\Big{)},$
@@ 5.2. Training and Inference
5.3. Selecting the Right Feature for Conditioning
@@
Selecting an appropriate context feature is one of the most critical and challenging aspects of contextual anomaly detection. The chosen context fundamentally shapes how the model interprets patterns in the data and can have a substantial impact on detection performance. As shown in Tables 3, 4, and 5, contextual models consistently outperform their non-contextual counterparts when the context is well chosen, underscoring its central role in effective anomaly detection. Achieving strong results with CWAE therefore requires a principled and effective strategy for selecting a context feature that is likely to yield good performance.
@@
We propose to frame context selection as a bilevel optimization problem (Sinha et al., 2018), where the goal is to optimize for the joint data likelihood $P(\mathbf{Y},\mathbf{C})=P(\mathbf{Y}\mid\mathbf{C})\,P(\mathbf{C})$. In this setting, both the model parameters and the context feature are treated as variables to optimize. The inner loop learns the model weights for a given context, while the outer loop searches for the context feature that yields the best generalization performance.
Naively training each conditional model to convergence is accurate, but can be prohibitively expensive. In practice, we only train the model for one epoch, and evaluate the model with validation loss afterwards. We observe that this approach serves as an effective proxy for generalization quality. Numerous prior works have demonstrated the effectiveness of leveraging an early-trained model for other tasks. Similar ideas have been applied for neural architecture search (Liu et al., 2019; Bender et al., 2018), neural network model pruning (You et al., 2020), dataset pruning (Paul et al., 2021) and learning with noisy labels (Liu et al., 2020; Zhang and Sabuncu, 2018). Prior works have also justified both empirically and theoretically the importance of the first few epochs for generalization of neural networks (Achille et al., 2017; Zhang et al., 2016; Frankle et al., 2020).
@@
To ensure consistency across models with different probabilistic factorizations—e.g., $P(\mathbf{Y}\mid\mathbf{C})$ vs. $P(\mathbf{Y},\mathbf{C})$—we evaluate all models using their joint log-likelihood. For conditional models, we apply the identity $P(\mathbf{Y},\mathbf{C})=P(\mathbf{Y}\mid\mathbf{C})\,P(\mathbf{C})$, where $P(\mathbf{C})$ is estimated empirically from the training data. This formulation also penalizes rare context values via the $\log P(\mathbf{C})$ term, favoring parsimonious and meaningful conditioning. For simplicity, we only condition on one feature at a time as the context feature. As such, $P(\mathbf{C})$ can be easily obtained from the training data.
Putting everything together we present the proposed procedure for selecting context features in Algorithm 1. This data-driven selection method avoids heuristic filtering and offers a principled framework for identifying useful context features.
@@
Figure 2. Validation loss curves for CWAE on the Census dataset, conditioned on different context features.
@@
Figure 3. AUCROC improvement achieved by applying Algorithm 1 to select context for CWAE across datasets.
5.4. Advantages of the Wasserstein Approach
@@
As outlined in Section 4, effective conditional modeling of $P(\mathbf{Y}\mid\mathbf{C})$ benefits from architectures that are both stable and efficient at inference. While Variational Autoencoders (VAEs) (Sohn et al., 2015) have been widely used for generative modeling, their non-deterministic encoders require multiple latent samples at inference to obtain stable anomaly scores, increasing computational cost and introducing score variability.
@@
Wasserstein Autoencoders (WAEs) replace the KL-divergence regularization in VAEs with an optimal transport–based term such as Maximum Mean Discrepancy (MMD), enabling fully deterministic encoders without the overfitting issues of unregularized autoencoders. Deterministic encoding removes the need for Monte Carlo sampling, making scoring faster and more stable—aligning with the robustness and efficiency goals discussed in Section 4.
@@
In this work, we instantiate CWAE with a deterministic encoder–decoder pair. This design choice reduces inference latency, improves score consistency, and maintains a well-regularized latent space, providing practical benefits over the Conditional VAE (CVAE) baseline while preserving the ability to model complex conditional distributions.
6. Experiments
6.1. Experiment Setup
6.1.1. Dataset Selection
@@
To evaluate contextual learning for anomaly detection, we selected datasets spanning diverse domains, including finance, cybersecurity, demographics, and network intrusion detection. Each dataset is designed to facilitate the study of anomaly detection with contextual learning, where anomalies are defined by specific criteria such as customer behavior, malicious events, or demographic thresholds.
@@
We started with datasets introduced in the AD literature, such as those discussed in (Han et al., 2022), and expanded our selection to include additional datasets originally used for multi-class classification. Collectively, we have compiled a suite of eight datasets for comprehensive empirical evaluation, which includes the Bank Marketing dataset (bank) (Moro et al., 2012), the Beth Cybersecurity dataset (beth) (Highnam et al., 2021), the Census dataset (census) (noa, 2000b), the CMC dataset (cmc) (Tjen-Sien Lim, 1999), the KDD dataset (kdd) (noa, 1999a), the LANL dataset (lanl) (Turcotte et al., 2019), the Solar Flares dataset (sf) (Unknown, 1989) and the Spotify dataset (spotify) (Choksi, 2021). Details about the datasets can be found in Table 1. These datasets vary significantly in size, feature count, and anomaly prevalence, providing a comprehensive foundation for robust evaluation.
6.1.2. Data Complexity Ranking
@@
Understanding dataset complexity is essential for ensuring that contextual learning methods are evaluated across a broad spectrum of dataset difficulties. By considering complexity, we can better assess how contextual learning behaves in both simpler and more challenging anomaly detection scenarios. This perspective allows us to examine whether the benefits of contextual learning hold consistently across varying levels of difficulty. To capture these complexities, we adopt several metrics from AD literature (Pang et al., 2021a), tailored to reflect the specific characteristics of our datasets. These metrics are defined as:
@@
Table 1. Overview of datasets with descriptive statistics.
@@
Table 2. Dataset complexity metrics with raw scores (Score), scaled scores (Scaled), and ranks (Rank) reported for each. Higher scores indicate greater complexity, while lower ranks correspond to higher complexity.
@@
6.1.3. Contextual Thresholding for AUCROC Calculation
@@
6.1.4. Benchmark Models
To effectively benchmark our proposed model, we selected six SOTA anomaly detection models from the DeepOD package (Xu, 2024): DSVDD (Ruff et al., 2018), RDP (Wang et al., 2019a), RCA (Liu et al., 2021), ICL (Shenkar and Wolf, 2022), DIF (Xu et al., 2023a), and SLAD (Xu et al., 2023b). Based on our research, these models represent leading techniques in the anomaly detection literature. The DeepOD package provides comprehensive implementations of these methods within a unified testbed, enabling consistent evaluation. Each model was tested using the default hyperparameter settings provided by DeepOD, ensuring comparability across experiments. All models were implemented without incorporating contextual considerations for anomaly detection.
In addition to the DeepOD-based models, we include the DTE (Livernoche et al., 2023) model, implemented from the available paper details. As one of the most recent diffusion-based approaches to anomaly detection, DTE introduces a fundamentally different architecture that broadens the diversity of our baseline set. We also evaluate the WAE (Tolstikhin et al., 2019) model as a non-contextual baseline. WAE shares the same architecture as our proposed CWAE but treats all features uniformly as content, without distinguishing contextual variables, making it a strong reference point for isolating the effects of contextual modeling.
6.2. Performance Evaluation
@@
Table 3. AUCROC scores, scaled complexity scores, and the selected context feature for all models across datasets. Bold and underlined values indicate the best AUCROC scores.
@@
Table 4. Model ranks for all datasets. Bold and underlined values indicate the best (lowest) rank.
@@
7. Ablation Studies
@@
Table 5. Comparison of AUCROC scores and ranks for WAE, CWAE, CWAE Best, and the best SOTA model per dataset. Bold and underlined values indicate the best AUCROC and best (lowest) rank.
7.1. Benefits of Contextual Thresholding
Figure 4. Thresholds for CWAE on the bank dataset, conditioned on different context features. Each point represents a threshold learned under a specific context group.
7.2. Upper Bound Performance with Optimal Context
@@
7.3. Quantifying the Value of Context
@@
  • Conditional Variational Autoencoder (CVAE) (Pol et al., 2020), which models the conditional distribution $P_{\theta}(\mathbf{Y}\mid\mathbf{C})$ of observations given contextual variables using a deep generative model with learned feature-wise reconstruction variance.
  • Conditional Mixture Density Network (CMDN) (Dai et al., 2025), which learns a neural network–parameterized Gaussian mixture model to capture $P(\mathbf{Y}\mid\mathbf{C})$, enabling mixture parameters to adapt dynamically based on context.
  • Contextual Anomaly Detection using Isolation Forest (CADI) (Yepmo et al., 2024), which extends the isolation forest algorithm with density-aware splits to jointly detect anomalies and explain them relative to local contextual clusters.
All models are implemented with only minor modifications from their original formulations, using the same embedding layers outlined in Section 5.1. For each model–dataset pair, we systematically iterated over all candidate features as potential context, designating exactly one feature as context in each run and treating the remainder as content. Performance was compared to a no-context baseline in which all features are treated as content. This controlled setting enables a direct, interpretable measure of how both context and model architecture contribute to performance improvements.

### v2

@@ 5.1.3 Loss Function
$\quad+\lambda\cdot\text{MMD}\Big(P(z),Q(z\mid X=\mathbf{Y},\mathbf{C})\Big),$ (1)
@@ 5.2 Training and Inference
5.3 Context Feature Selection
@@
Selecting an appropriate context feature is one of the most critical and challenging aspects of the contextual learning framwork. The chosen context fundamentally shapes how the model interprets patterns in the data and can have a substantial impact on detection performance. As shown in Tables 3, 4, and 5, contextual models consistently outperform their non-contextual counterparts when the context is well chosen, underscoring its central role in effective anomaly detection. Achieving strong results therefore requires a principled and effective strategy for selecting a context feature that is likely to yield good performance.
@@
We propose to frame context selection as a bilevel optimization problem sinha2018overview, where the goal is to optimize for the joint data likelihood $P(\mathbf{Y},\mathbf{C})=P(\mathbf{Y}\mid\mathbf{C})\,P(\mathbf{C})$. In this setting, both the model parameters and the context feature are treated as variables to optimize. The inner loop learns the model weights for a given context, while the outer loop searches for the context feature that yields the best generalization performance.
To enable reliable context selection, our approach uses a small validation subset drawn from data that is expected to be mostly normal. This requirement is minimal, since only enough samples are needed to compute stable loss estimates, not labeled anomalies. Such assumptions are common in practical anomaly detection settings, where a short baseline period or trusted historical data is usually available (e.g., initial system deployment, known-good historical data). The validation data is used only to evaluate generalization performance for context selection, not to train the anomaly detector itself. The framework remains fully unsupervised with respect to anomaly labels, as no labeled anomalies are used at any stage. When a clean validation subset is unavailable, practitioners can instead reserve a small portion of the training data as a proxy validation set, although this may slightly reduce selection reliability.
Naively training each conditional model to convergence is accurate, but can be prohibitively expensive. In practice, we only train the model for one epoch, and evaluate the model with validation loss afterwards.
Our single-epoch evaluation strategy is motivated by extensive empirical and theoretical evidence showing that early training dynamics are strongly predictive of final generalization performance. Numerous prior works have successfully leveraged early-trained models for neural architecture search liu2019darts; bender2018understanding, model pruning you2020drawing, dataset pruning paul2021deep, and learning with noisy labels liu2020early; zhang2018generalized. Studies have also shown that the first few epochs establish key characteristics of the learned representation achille2017critical; zhang2016understanding; frankle2020early. In our context selection setting, the goal is not absolute performance but rather the relative ranking among context features, a signal that emerges early in training. Figure 3 supports this empirically, as the early-epoch proxy achieves better performance on five of eight datasets.
@@
To ensure consistency across models with different probabilistic factorizations—e.g., $P(\mathbf{Y}\mid\mathbf{C})$ vs. $P(\mathbf{Y},\mathbf{C})$—we evaluate all models using their joint log-likelihood. For conditional models, we apply the identity $P(\mathbf{Y},\mathbf{C})=P(\mathbf{Y}\mid\mathbf{C})\,P(\mathbf{C})$, where $P(\mathbf{C})$ is estimated empirically from the training data. This formulation also penalizes rare context values via the $\log P(\mathbf{C})$ term, favoring parsimonious and meaningful conditioning. As motivated in Section 3.1, we only condition on one feature at a time as the context feature. As such, $P(\mathbf{C})$ can be easily obtained from the training data.
@@
Putting everything together we present the proposed procedure for selecting context features in Algorithm 1. This data-driven selection method avoids heuristic filtering and provides a principled mechanism for identifying informative context features within the contextual learning framework.
@@
Figure 2: Validation loss curves for CWAE on the Census dataset, conditioned on different context features.
@@
Figure 3: AUCROC improvement achieved by applying Algorithm 1 to select context for CWAE across datasets.
5.4 Advantages of the Wasserstein Approach
@@
As outlined in Section 4, effective conditional modeling of $P(\mathbf{Y}\mid\mathbf{C})$ benefits from architectures that are both stable and efficient at inference. While Variational Autoencoders (VAEs) sohn2015cvae have been widely used for generative modeling, their non-deterministic encoders require multiple latent samples at inference to obtain stable anomaly scores, increasing computational cost and introducing score variability.
@@
Wasserstein Autoencoders (WAEs) replace the KL-divergence regularization in VAEs with an optimal transport-based term such as Maximum Mean Discrepancy (MMD), enabling fully deterministic encoders without the overfitting issues of unregularized autoencoders. Deterministic encoding removes the need for Monte Carlo sampling, making scoring faster and more stable—aligning with the robustness and efficiency goals discussed in Section 4.
@@
In this work, we instantiate CWAE with a deterministic encoder-decoder pair. This design choice reduces inference latency, improves score consistency, and maintains a well-regularized latent space, providing practical benefits over the Conditional VAE (CVAE) baseline while preserving the ability to model complex conditional distributions.
6 Experiments
6.1 Experiment Setup
6.1.1 Dataset Selection
@@
To evaluate the effectiveness of the contextual learning framework, we instantiate it using CWAE and assess its performance on selected datasets spanning diverse domains, including finance, cybersecurity, demographics, and network intrusion detection. Each dataset is designed to facilitate the study of anomaly detection with contextual learning, where anomalies are defined by specific criteria such as customer behavior, malicious events, or demographic thresholds.
@@
We started with datasets introduced in the AD literature, such as those discussed in han_adbench_2022, and expanded our selection to include additional datasets originally used for multi-class classification. Collectively, we have compiled a suite of eight datasets for comprehensive empirical evaluation, which includes the Bank Marketing dataset (bank) moro_bank_2012, the Beth Cybersecurity dataset (beth) highnam_beth_nodate, the Census dataset (census) noauthor_census-income_2000, the CMC dataset (cmc) tjen-sien_lim_contraceptive_1999, the KDD dataset (kdd) noauthor_kdd_1999, the LANL dataset (lanl) turcotte_unified_2018, the Solar Flares dataset (sf) unknown_solar_1989 and the Spotify dataset (spotify) spotify_ds. Details about the datasets can be found in Table 1. These datasets vary significantly in size, feature count, and anomaly prevalence, providing a comprehensive foundation for robust evaluation. Complete implementation and evaluation details to facilitate reproducibility are provided in the Appendix A.
6.1.2 Data Complexity Ranking
@@
Understanding dataset complexity is essential for ensuring that contextual learning methods are evaluated across a broad spectrum of dataset difficulties. By considering complexity, we can better assess how contextual learning behaves in both simpler and more challenging anomaly detection scenarios. This perspective allows us to examine whether the benefits of contextual learning hold consistently across varying levels of difficulty. To capture these complexities, we adopt several metrics from AD literature pang_homophily_2021, tailored to reflect the specific characteristics of our datasets. These metrics are defined as:
@@
Table 1: Overview of datasets with descriptive statistics.
@@
Table 2: Dataset complexity metrics with raw scores (Score), scaled scores (Scaled), and ranks (Rank) reported for each. Higher scores indicate greater complexity, while lower ranks correspond to higher complexity.
@@
6.1.3 Contextual Thresholding for AUCROC Calculation
@@
6.1.4 Benchmark Models
To effectively benchmark our proposed model, we selected six SOTA anomaly detection models from the DeepOD package xu_xuhongzuodeepod_2024: DSVDD ruff_deep_2018, RDP wang_unsupervised_2020, RCA liu_rca_2021, ICL shenkar_anomaly_2022, DIF xu_deep_2023, and SLAD xu_fascinating_2023. Based on our research, these models represent leading techniques in the anomaly detection literature. The DeepOD package provides comprehensive implementations of these methods within a unified testbed, enabling consistent evaluation. Each model was tested using the default hyperparameter settings provided by DeepOD, ensuring comparability across experiments. All models were implemented without incorporating contextual considerations for anomaly detection.
In addition to the DeepOD-based models, we include the DTE livernoche2023diffusion model, implemented from the available paper details. As one of the most recent diffusion-based approaches to anomaly detection, DTE introduces a fundamentally different architecture that broadens the diversity of our baseline set. We also evaluate the WAE tolstikhin_wasserstein_2019 model as a non-contextual baseline. WAE shares the same architecture as our proposed CWAE but treats all features uniformly as content, without distinguishing contextual variables, making it a strong reference point for isolating the effects of contextual modeling.
6.2 Performance Evaluation
@@
Table 3: AUCROC scores, scaled complexity scores, and the selected context feature for all models across datasets. Bold and underlined values indicate the best AUCROC scores.
@@
Table 4: Model ranks for all datasets. Bold and underlined values indicate the best (lowest) rank.
@@
7 Ablation Studies
7.1 Benefits of Contextual Thresholding
Figure 4: Thresholds for CWAE on the bank dataset, conditioned on different context features. Each point represents a threshold learned under a specific context group.
7.2 Upper Bound Performance with Optimal Context
@@
Table 5: Comparison of AUCROC scores and ranks for WAE, CWAE, CWAE Best, and the best SOTA model per dataset. Bold and underlined values indicate the best AUCROC and best (lowest) rank.
@@
7.3 Quantifying the Value of Context
@@
  • Conditional Variational Autoencoder (CVAE) pol2020anomaly, which models the conditional distribution $P_{\theta}(\mathbf{Y}\mid\mathbf{C})$ of observations given contextual variables using a deep generative model with learned feature-wise reconstruction variance.
  • Conditional Mixture Density Network (CMDN) dai2025deep, which learns a neural network-parameterized Gaussian mixture model to capture $P(\mathbf{Y}\mid\mathbf{C})$, enabling mixture parameters to adapt dynamically based on context.
  • Contextual Anomaly Detection using Isolation Forest (CADI) yepmo2024cadi, which extends the isolation forest algorithm with density-aware splits to jointly detect anomalies and explain them relative to local contextual clusters.
All models are implemented with only minor modifications from their original formulations, using the same embedding layers outlined in Section 5.1 (see Appendix A for more details). For each model-dataset pair, we systematically iterated over all candidate features as potential context, designating exactly one feature as context in each run and treating the remainder as content. Performance was compared to a no-context baseline in which all features are treated as content. This controlled setting enables a direct, interpretable measure of how both context and model architecture contribute to performance improvements.